# On the multiplier spectrum of polynomials

### v1

@@ 4.1. Marked critical points and a theorem of Favre-Gauthier
On the parameter space ${\rm Poly}^{d}$ (or the moduli space ${\rm MPoly}^{d}$), the critical points do not form global morphisms. This causes difficulties when working with information about critical points (e.g., PCF maps and stable preperiodic critical points). To address this, it is necessary to consider critically marked polynomials via a base change of the parameter space (or the moduli space). In this subsection, we gather classical constructions of critically marked polynomials; see [FG22, §2.1].
A critically marked polynomial of degree $d\geq 2$ is a tuple $(P,c_{0},\dots,c_{d-2})$ where $P\in{\rm Poly}^{d}$ and $c_{0},\dots,c_{d-2}$ are the critical points (other than $\infty$) of $P$, counted with multiplicity. Let ${\rm MPcrit}^{d}$ denote the quotient of the space of critically marked polynomials of degree $d$ by the natural action of the group $\rm Aff$} of affine transformations. There is a canonical forgetful morphism ${\rm MPcrit}^{d}\to{\rm MPoly}^{d}$.
@@
$\eta:\mathbb{A}^{d-1}\to{\rm MPcrit}^{d},(c,a)=(c_{1},\dots,c_{d-2},a)\mapsto[P_{c,a}(z)],$
@@
and $\sigma_{j}(c)$ is the $j$-th elementary symmetric polynomial in $(c_{1},\dots,c_{d-2})$ for $1\leq j\leq d-2$. Note that the (finite) critical points of $P_{c,a}$ are exactly $c_{0}:=0,c_{1},\dots,c_{d-2}$. The morphism $\eta:\mathbb{A}^{d-1}\to{\rm MPcrit}^{d}$ is defined over $\mathbb{Q}$ and is d(d1)d(d-1)-to-one, so $\mathbb{A}^{d-1}$ is a finite ramified cover of ${\rm MPcrit}^{d}$.
The following theorem is a part of [FG22, Theorem B] and will be needed to prove Theorem 1.9.
Theorem 4.1 (Favre-Gauthier).
Let $(P,a)$ and $(Q,b)$ be non-exceptional active dynamical pairs parametrized by an irreducible algebraic curve $C$, of respective degrees $d,\delta\geq 2$, defined over a number field $K$. Then the following are equivalent:
@@
    The set $\{t\in C(\overline{K}):a(t)\in{\rm PrePer}(P(t))\text{ and }b(t)\in{\rm PrePer}(Q(t))\}$ is infinite;
@@
    There exist integers $N,M\geq 1$, $r,s\geq 0$, and families $R,\pi,\tau$ of polynomials of degree $\geq 1$ parametrized by $C$ such that τP^N=Rτ and πQ^M=Rπ with $\tau(P^{\circ N}(a))=\pi(Q^{\circ M}(b))$.
Here, a dynamical pair parametrized by $C$ of degree $d$ is a pair $(P,a)$ where $P:C\to{\rm MPoly}^{d}$ is a family of polynomials of degree $d$ over $C$, and $a:C\to\mathbb{A}^{1}$ is a morphism (i.e., a marked point on the family $P$). The dynamical pair $(P,a)$ is called active (or not stably preperiodic) if for all distinct integers $n,m\geq 0$ we have $(P(t))^{\circ n}(a(t))\neq(P(t))^{\circ m}(a(t))$ on $C$. It is called non-exceptional if the polynomial $P(t)$ is not exceptional for every $t\in U(\overline{K})$, where $U$ is a suitable dense open subset of $C$. If two pairs $(P,a)$ and $(Q,b)$ satisfy the property in (1), then they are called entangled.
@@ 4.2. Description of the non-injective locus
We now prove Theorem 1.9, relying on results of Pakovich and Theorem 4.1 of Favre-Gauthier.
From the proof of [JX25, Lemma 3.5], we obtain the following elementary lemma:
@@
Let $f$ be a polynomial map of degree $d\geq 2$. Then its multiplier spectrum $(S_{n}(f))_{n=1}^{\infty}$ determines the number ${\rm sac}(f)\in\{0,1,\dots,2d-2\}$ of superattracting cycles of $f$ in $\mathbb{C}$ (counted without multiplicity). (The bound ${\rm sac}(f)\leq 2d-2$ follows from the Riemann–Hurwitz formula.)
@@
Let $\phi:C\to{\rm MPoly}^{d}$ be a non-isotrivial family of polynomials of degree $d\geq 2$, parametrized by an irreducible curve $C$. Assume that $\phi$ factors through $\mathbb{A}^{d-1}\xrightarrow{\eta}{\rm MPcrit}^{d}\to{\rm MPoly}^{d}$. Let $m(\phi)\geq 0$ be the minimal nonnegative integer such that the subset $\{t\in C(\mathbb{C}):{\rm sac}(\phi(t))=m\}$ is infinite. Define
$Z(\phi):=\{t\in C(\mathbb{C}):{\rm sac}(\phi(t))>m(\phi)\},$
and let $Z^{\prime}(\phi)$ be the set of all $t\in C(\mathbb{C})$ such that $c_{j}$ is not a stably preperiodic marked point of $\phi$ but $c_{j}(t)\in{\rm Per}(\phi(t))$ for some $0\leq j\leq d-1$.
@@
By Lemma 4.2, the number $m(t)$ depends only on the multiplier spectrum, and hence so does $Z(\phi)$. Note that $m(\phi)$ does not change after removing finitely many points of $C(\mathbb{C})$. In the following argument, we are free to remove finitely many points of $C(\mathbb{C})$.
Let $0\leq i,j\leq d-1$ be indices such that both $c_{i}$ and $c_{j}$ are stably periodic for $\phi$. Assume that $n,m\in\mathbb{Z}_{>0}$ are minimal such that $\phi(t)^{\circ n}(c_{i}(t))=c_{i}(t)$ and $\phi(t)^{\circ m}(c_{j}(t))=c_{j}(t)$ on all of $C$. For every integer $0\leq r\leq n$, the subset
$\{t\in C(\mathbb{C}):c_{j}(t)=\phi(t)^{\circ r}(c_{i}(t))\}$
@@
$\mathcal{O}_{\phi(t)}(c_{i}(t))=\mathcal{O}_{\phi(t)}(c_{j}(t))\quad\text{for all }t\in C(t),$
@@
$\mathcal{O}_{\phi(t)}(c_{i}(t))\cap\mathcal{O}_{\phi(t)}(c_{j}(t))=\emptyset\quad\text{for all }t\in C(t).$
In the first case, write $i\sim j$. Let
$P(\phi):=\{0\leq i\leq d-1:\ c_{i}\text{ is stably periodic for }\phi\}.$
@@
Let $0\leq i\leq d-1$ be such that $c_{i}$ is stably preperiodic but not stably periodic for $\phi$. Assume $k>l\geq 0$ are minimal such that $\phi(t)^{\circ k}(c_{i}(t))=\phi(t)^{\circ l}(c_{i}(t))$ on all of $C$. Then $l>0$. For every $l\leq r\leq k$, the subset
$\{t\in C(\mathbb{C}):c_{i}(t)=\phi(t)^{\circ r}(c_{i}(t))\}$
is finite, since $c_{i}$ is not stably periodic. After removing finitely many points of $C(\mathbb{C})$, we may assume that
(4.1) $c_{i}(t)\notin{\rm Per}(\phi(t))\quad\text{for all }t\in C(\mathbb{C}).$
It is well known that PCF parameters do not form families in ${\rm MPoly}^{d}$. Thus,
(4.2) $A(\phi):=\{0\leq i\leq d-1:\ c_{i}\text{ is not stably preperiodic for }\phi\}\neq\emptyset.$
Fix any $i\in A(\phi)$. Similarly (after removing finitely many points of $C(\mathbb{C})$), we may assume that $c_{i}(t)\notin\mathcal{O}_{\phi(t)}(c_{k}(t))$ for all $t\in C(\mathbb{C})$ and all $k\in P(\phi)$. By [JX25, Lemma 2.4],
(4.3) $\{t\in C(\mathbb{C}):c_{i}(t)\in{\rm Per}(\phi(t))\}\ \text{is infinite.}$
On the other hand, since $c_{i}$ is not stably preperiodic for $\phi$, the set
$\{t\in C(\mathbb{C}):\phi(t)^{\circ n}(c_{i}(t))=c_{i}(t)\}$
is finite for every $n\geq 1$. Thus, by the uncountability of $C(\mathbb{C})$,
(4.4) $\bigcap_{k\in A(\phi)}\{t\in C(\mathbb{C}):c_{k}(t)\notin{\rm Per}(\phi(t))\}=+\infty.$
From the case-by-case analysis above, it is easy to see that
@@
Indeed, we have ${\rm sac}(\phi(t))\geq\#(P(\phi)/\sim)$ for all $t\in C(\mathbb{C})$ by the analysis of stably periodic $c_{i}$’s, hence $m(\phi)\geq\#(P(\phi)/\sim)$. From (4.1) and (4.4), there are uncountably many $t\in C(\mathbb{C})$ with ${\rm sac}(\phi(t))=\#(P(\phi)/\sim)$. Hence $m(\phi)=\#(P(\phi)/\sim)$.
By (4.2) and (4.3), $Z^{\prime}(\phi)$ is infinite. For every $t\in Z^{\prime}(\phi)$,
${\rm sac}(\phi(t))\geq\#(P(\phi)/\sim)+1>m(\phi),$
so $Z^{\prime}(\phi)\subseteq Z(\phi)$; hence $Z(\phi)$ is also infinite. Conversely, for every $t\in C(\mathbb{C})\setminus Z^{\prime}(\phi)$, by (4.1) we have ${\rm sac}(\phi(t))=\#(P(\phi)/\sim)=m(\phi)$, so $Z(\phi)\subseteq Z^{\prime}(\phi)$. Therefore $Z(\phi)=Z^{\prime}(\phi)$, completing the proof. ∎
@@
After a suitable base change of $C$ (restricting to a suitable non-empty affine open subset), we may assume that $\phi_{1}(t)$ and $\phi_{2}(t)$ are non-exceptional such that $(\phi_{1}(t),\phi_{2}(t))\in{\rm PNI}_{d}$ for every $t\in C(\mathbb{C})$, and that $\phi_{1},\phi_{2}$ factor through
$\mathbb{A}^{d-1}\xrightarrow{\eta}{\rm MPcrit}^{d}\to{\rm MPoly}^{d}.$
We continue to denote by $\phi_{1},\phi_{2}$ the corresponding morphisms $C\to\mathbb{A}^{d-1}$.
By Proposition 4.3, we have $m(\phi_{1})=m(\phi_{2})$, $Z(\phi_{1})=Z(\phi_{2})$, and the subsets $Z^{\prime}(\phi_{1})$ and $Z^{\prime}(\phi_{2})$ are infinite with finite symmetric difference $Z^{\prime}(\phi_{1})\Delta Z^{\prime}(\phi_{2})$. For $i=1,2$ and $k\in A(\phi_{i})$, set
$Z_{ik}=\{t\in C(\mathbb{C}):c_{k}(t)\in{\rm Per}(\phi_{i}(t))\}.$
@@
Note that $Z_{ik}=\{t\in C(\overline{\mathbb{Q}}):c_{k}(t)\in{\rm Per}(\phi_{i}(t))\}$ since all the objects are defined over $\overline{\mathbb{Q}}$. Then $Z^{\prime}(\phi_{i})=\bigcup_{k\in A(\phi_{i})}Z_{ik}$ for $i=1,2$. Since $Z^{\prime}(\phi_{1})\cap Z^{\prime}(\phi_{2})$ is infinite, we can choose $k\in A(\phi_{1})$ and $l\in A(\phi_{2})$ such that $(Z_{1k}\cap Z_{2l})=+\infty$. After renumbering, assume $k=l=0$. Then 0 is not stably preperiodic for $\phi_{1}$ and $\phi_{2}$, and we obtain two active dynamical pairs $(\phi_{1},0)$ and $(\phi_{2},0)$ with
$\{t\in C(\mathbb{C}):0\in{\rm Per}(\phi_{1}(t))\cap{\rm Per}(\phi_{2}(t))\}=+\infty$
(since $(Z_{10}\cap Z_{20})=+\infty$). In the sense of Favre–Gauthier [FG22], the pairs $(\phi_{1},0)$ and $(\phi_{2},0)$ are entangled.
@@
Applying Theorem 4.1, there exist $N,M\in\mathbb{Z}_{>0}$ and families $R,\tau,\pi$ of polynomials (of degree $\geq 1$) parametrized by $C$ such that
$\tau\circ\phi_{1}^{\circ N}=R\circ\tau,\qquad\pi\circ\phi_{2}^{\circ M}=R\circ\pi,\qquad\tau(\phi_{1}^{\circ N}(0))=\pi(\phi_{2}^{\circ M}(0)).$
Note that $d^{N}=\deg(\phi_{1}^{\circ N})=\deg(R)=\deg(\phi_{2}^{\circ M})=d^{M}$, hence $N=M$. By composing with $\phi_{1},\phi_{2}$, respectively, we may assume that the families $\tau$ and $\pi$ are of degree at least $2$.
@@
Fix $t\in C(\mathbb{C})$. Write $h_{1}=\phi_{1}(t)^{\circ N}$ and $h_{2}=\phi_{2}(t)^{\circ N}$. For simplicity, we write $R,\tau,\pi$ for $R(t),\tau(t),\pi(t)$, respectively. Then
$\tau\circ h_{1}=R\circ\tau,\quad\pi\circ h_{2}=R\circ\pi,\quad\tau(h_{1}(0))=\pi(h_{2}(0)).$
@@
Assume $h_{1}$ and $h_{2}$ are not equivalent. We will show that there exist positive integers $k_{1},k_{2},l$, a non-constant polynomial $V(z)\in\mathbb{C}[z]\setminus\mathbb{C}$, and $\zeta\in\mathbb{C}$, with
$2\leq k:={\rm lcm}(k_{1},k_{2})\leq d^{N},\quad\gcd(d,k)=1,\quad V(0)\neq 0,\quad\zeta^{k_{2}}=1,$
@@
where $k_{j}^{\prime}=k/k_{j}$ for $j=1,2$.
Since $h_{1}\not\sim h_{2}$, there exists $j\in\{1,2\}$ such that $h_{j}\not\sim R$. Without loss of generality, assume $h_{1}\not\sim R$. By Theorem 2.10, there exist non-constant rational maps $\tau_{0},\tau_{1},\tilde{h}_{1}$ satisfying:
  • $R$ is a generalized Lattès map;
  • $\tau=\tau_{0}\circ\tau_{1}$ with $\deg(\tau_{0})\geq 2$;
@@
  • $R:\mathcal{O}_{2}^{\tau_{0}}\to\mathcal{O}_{2}^{\tau_{0}}$ and $\tilde{h}_{1}:\mathcal{O}_{1}^{\tau_{0}}\to\mathcal{O}_{1}^{\tau_{0}}$ are minimal holomorphic maps between orbifolds;

### v2

@@ 4.1. Marked critical points and a theorem of Favre-Gauthier
On the parameter space $\mathrm{Poly}^{d}$ (or the moduli space $\mathrm{MPoly}^{d}$), the critical points are not given by global morphisms. This causes difficulties when working with information about critical points (e.g., PCF maps and persistently preperiodic critical points). To address this, it is necessary to consider critically marked polynomials via a base change of the parameter space (or the moduli space). In this subsection, we review classical constructions of critically marked polynomials; see [FG22, § 2.1] and [DF08, § 5].
A critically marked polynomial of degree $d\geq 2$ is a tuple $(P,c_{0},\dots,c_{d-2})$ where $P\in\mathrm{Poly}^{d}$ and $c_{0},\dots,c_{d-2}$ are the critical points (other than $\infty$) of $P$, counted with multiplicity. The space of critically marked polynomials of degree $d\geq 2$ is the closed subvariety $\mathrm{Pcrit}^{d}$ of $\mathrm{Poly}^{d}\times\mathbb{A}^{d-1}$ given by
$\left\{(P=\sum_{i=0}^{d}a_{i}z^{i},c_{0},c_{1},\dots,c_{d-2})\in\mathrm{Poly}^{d}\times\mathbb{A}^{d-1}\colon\sum_{j=1}^{d}ja_{j}z^{j-1}=da_{d}\prod_{k=0}^{d-2}(z-c_{k})\right\}.$
Let $\mathrm{MPcrit}^{d}$ denote the quotient $\mathrm{Pcrit}^{d}/\mathrm{Aff}$, where the action is given by
$\sigma\cdot\left(P,c_{0},\dots,c_{d-2}\right)=\left(\sigma\circ P\circ\sigma^{-1},\sigma(c_{0}),\dots,\sigma(c_{d-2})\right),$
for $\sigma\in\mathrm{Aff}$ and $\left(P,c_{0},\dots,c_{d-2}\right)\in\mathrm{Pcrit}^{d}$. There is a canonical forgetful morphism $\mathrm{MPcrit}^{d}\to\mathrm{MPoly}^{d}$.
@@
$\eta:\mathbb{A}^{d-1}\to\mathrm{MPcrit}^{d},(c,a)=(c_{1},\dots,c_{d-2},a)\mapsto\left[\left(P_{c,a}(z),0,c_{1},\dots,c_{d-2}\right)\right],$
@@
and $\sigma_{j}(c)$ is the $j$-th elementary symmetric polynomial in $(c_{1},\dots,c_{d-2})$ for $1\leq j\leq d-2$. Note that the (finite) critical points of $P_{c,a}$ are exactly $c_{0}:=0,c_{1},\dots,c_{d-2}$. The morphism $\eta:\mathbb{A}^{d-1}\to\mathrm{MPcrit}^{d}$ is defined over $\mathbb{Q}$ and is a finite ramified covering of degree d(d1)d(d-1) (see [DF08, Proposition 5.1]). Let $\eta_{0}:\mathbb{A}^{d-1}\to\mathrm{Poly}^{d}$ be the morphism defined by $\eta_{0}(c,a)=P_{c,a}$.
The following theorem is a part of [FG22, Theorem B] and is essential for the proof of Theorem 1.10.
Theorem 4.1 (Favre and Gauthier).
Let $(P,a)$ and $(Q,b)$ be non-exceptional active dynamical pairs parametrized by an irreducible algebraic curve $C$, of respective degrees $d,\delta\geq 2$, defined over a number field $K$. Then the following conditions are equivalent:
@@
    The set $\{t\in C(\overline{\mathbb{Q}})\colon a(t)\in\mathrm{PrePer}(P(t))\text{ and }b(t)\in\mathrm{PrePer}(Q(t))\}$ is infinite;
@@
    There exist integers $N,M\geq 1$ and families $R,\pi,\tau$ of polynomials of degree $\geq 1$ parametrized by $C$ such that τP^N=Rτ and πQ^M=Rπ.
Here, a dynamical pair parametrized by $C$ of degree $d$ is a pair $(P,a)$ where $P:C\to\mathrm{Poly}^{d}$ is a family of polynomials of degree $d$ over $C$, and $a:C\to\mathbb{A}^{1}$ is a morphism (i.e., a marked point of the family $P$). The dynamical pair $(P,a)$ is called active (or not persistently preperiodic) if for all integers $n>m\geq 0$, we have $(P(t))^{\circ n}(a(t))\neq(P(t))^{\circ m}(a(t))$ in the function field of $C$. It is called non-exceptional if the polynomial $P(t)$ is non-exceptional for every $t\in U(\overline{\mathbb{Q}})$, where $U$ is a suitable non-empty Zariski open subset of $C$. If two pairs $(P,a)$ and $(Q,b)$ satisfy the property in (1), then they are called entangled.
@@ 4.2. Description of the non-injective locus
We prove Theorem 1.10, relying on results of Pakovich and Theorem 4.1 of Favre-Gauthier.
Following the proof of [JX25, Lemma 3.5], we obtain the following elementary lemma:
@@
Let $f$ be a polynomial map of degree $d\geq 2$. Then its multiplier spectrum $(S_{n}(f))_{n=1}^{\infty}$ determines the number $\mathrm{sac}(f)\in\{0,1,\dots,d-1\}$ of superattracting cycles of $f$ in $\mathbb{C}$ (counted without multiplicity).
Since $\infty$ is a totally ramified fixed point of $f$, the bound $\mathrm{sac}(f)\leq d-1$ follows from the Riemann–Hurwitz formula. Note that $\mathrm{sac}(\cdot)$ descends to the moduli space $\mathrm{MPoly}^{d}$.
@@
Let $\phi:C\to\mathrm{Poly}^{d}$ be a non-isotrivial family of polynomials of degree $d\geq 2$, parametrized by an irreducible curve $C$. Assume that $\phi$ factors through $\eta_{0}:\mathbb{A}^{d-1}\to\mathrm{Poly}^{d},(c,a)\mapsto P_{c,a}$ via some morphism $\psi:C\to\mathbb{A}^{d-1}$. Fix such a $\psi$. Let $m(\phi)\geq 0$ be the smallest nonnegative integer $m$ such that
$\{t\in C(\mathbb{C})\colon\mathrm{sac}(\phi(t))=m\}$
is infinite. Define
$Z(\phi):=\{t\in C(\mathbb{C})\colon\mathrm{sac}(\phi(t))>m(\phi)\},$
and let $Z^{\prime}(\phi)$ be the set of all $t\in C(\mathbb{C})$ such that $c_{j}\circ\psi$ is not a persistently $\phi$-preperiodic marked point but $c_{j}(\psi(t))\in\mathrm{Per}(\phi(t))$ for some $0\leq j\leq d-2$. Here $c_{j}$ is the coordinate map $\mathbb{A}^{d-1}\to\mathbb{A}^{1},(c,a)\mapsto c_{j}$ for $1\leq j\leq d-2$, and $c_{0}=0:\mathbb{A}^{d-1}\to\mathbb{A}^{1}$.
@@
By Lemma 4.2, the number $m(\phi)$ depends only on the multiplier spectrum, and hence so does $Z(\phi)$. Note that $m(\phi)$ does not change after removing finitely many points of $C(\mathbb{C})$. Hence it suffices to prove that $Z(\phi)=Z^{\prime}(\phi)$ is infinite, after removing finitely many points of $C(\mathbb{C})$.
For simplicity, we denote the marked point $c_{j}\circ\psi:C\to\mathbb{A}^{1}$ of $\phi$ by $a_{j}$, for $0\leq j\leq d-2$.
Let $(i,j)\in\{0,\dots,d-2\}^{2}$ be a pair of indices such that both $a_{i}$ and $a_{j}$ are persistently $\phi$-periodic. Let $n,m\in\mathbb{Z}_{>0}$ be minimal such that
$\phi(t)^{\circ n}(a_{i}(t))=a_{i}(t)\quad\text{and}\quad\phi(t)^{\circ m}(a_{j}(t))=a_{j}(t)$
in the function field of $C$. For every integer $0\leq r<n$, the subset
$\{t\in C(\mathbb{C})\colon a_{j}(t)=\phi(t)^{\circ r}(a_{i}(t))\}$
@@
(4.1) $O_{\phi(t)}(a_{i}(t))=O_{\phi(t)}(a_{j}(t))\quad\text{for all }t\in C(\mathbb{C}),$
@@
(4.2) $O_{\phi(t)}(a_{i}(t))\cap O_{\phi(t)}(a_{j}(t))=\emptyset\quad\text{for all }t\in C(\mathbb{C}).$
Here for $g\in\mathbb{C}(z)$ and $w\in\mathbb{P}^{1}(\mathbb{C})$, $O_{g}(w)$ denotes the forward $g$-orbit of $w$. In the first case (4.1), we write $i\sim j$. Let
$P(\phi):=\{0\leq i\leq d-2\colon a_{i}\text{ is persistently }\phi\text{-periodic}\}.$
@@
Let $i\in\{0,\dots,d-2\}$ be an index such that $a_{i}$ is persistently $\phi$-preperiodic but not persistently $\phi$-periodic. Let $k>l\geq 0$ be minimal integers such that
$\phi(t)^{\circ k}(a_{i}(t))=\phi(t)^{\circ l}(a_{i}(t))$
in the function field of $C$. Then $l>0$. For every integer $l\leq r<k$, the subset
$\{t\in C(\mathbb{C})\colon a_{i}(t)=\phi(t)^{\circ r}(a_{i}(t))\}$
is finite, since $a_{i}$ is not persistently periodic. After removing finitely many points of $C(\mathbb{C})$, for all such indices $i$, we may assume that
(4.3) $a_{i}(t)\notin\mathrm{Per}(\phi(t))\quad\text{for all }t\in C(\mathbb{C}).$
It is well-known that PCF parameters do not form families in $\mathrm{MPoly}^{d}$. Since $\phi$ is non-isotrivial, we conclude that
(4.4) $A(\phi):=\{0\leq i\leq d-2\colon a_{i}\text{ is not persistently }\phi\text{-preperiodic}\}\neq\emptyset.$
After removing finitely many points of $C(\mathbb{C})$ and arguing similarly, we may assume that
(4.5) $c_{i}(t)\notin O_{\phi(t)}(c_{k}(t))$
for all $(i,k)\in A(\phi)\times P(\phi)$ and all $t\in C(\mathbb{C})$. Fix any $i\in A(\phi)$. By [JX25, Lemma 2.4],
(4.6) $\{t\in C(\mathbb{C})\colon a_{i}(t)\in\mathrm{Per}(\phi(t))\}=+\infty.$
On the other hand, since $a_{i}$ is not persistently $\phi$-preperiodic, we have
$\{t\in C(\mathbb{C})\colon\phi(t)^{\circ n}(a_{i}(t))=a_{i}(t)\}<+\infty,$
for every $n\in\mathbb{Z}_{>0}$. So $\{t\in C(\mathbb{C})\colon a_{i}(t)\in\mathrm{Per}(\phi(t))\}$ is countable. Due to the uncountability of $C(\mathbb{C})$,
(4.7) $\bigcap_{k\in A(\phi)}\{t\in C(\mathbb{C})\colon a_{k}(t)\notin\mathrm{Per}(\phi(t))\}=+\infty.$
Based on the case analysis above, it is evident that
@@
Indeed, we have $\mathrm{sac}(\phi(t))\geq\#(P(\phi)/\sim)$ for all $t\in C(\mathbb{C})$ by the analysis of persistently $\phi$-periodic $a_{i}$’s, hence $m(\phi)\geq\#(P(\phi)/\sim)$. By (4.3) and (4.7), there are infinitely many (in fact, uncountably many) $t\in C(\mathbb{C})$ such that $\mathrm{sac}(\phi(t))=\#(P(\phi)/\sim)$. Hence $m(\phi)=\#(P(\phi)/\sim)$ by definition.
By (4.4) and (4.6), $Z^{\prime}(\phi)$ is infinite. For every $t\in Z^{\prime}(\phi)$, from (4.5) we see that
$\mathrm{sac}(\phi(t))\geq\#(P(\phi)/\sim)+1>m(\phi),$
so $Z^{\prime}(\phi)\subseteq Z(\phi)$; hence $Z(\phi)$ is also infinite. Conversely, for every $t\in C(\mathbb{C})\setminus Z^{\prime}(\phi)$, by (4.3) we have $\mathrm{sac}(\phi(t))=\#(P(\phi)/\sim)=m(\phi)$, so $Z(\phi)\subseteq Z^{\prime}(\phi)$. Therefore $Z(\phi)=Z^{\prime}(\phi)$, completing the proof. ∎
@@
After a suitable base change of $C$ (and restricting to a suitable non-empty affine open subset) over $\overline{\mathbb{Q}}$, we may assume that the following hold:
  • for every $t\in C(\mathbb{C})$, both $\phi_{1}(t)$ and $\phi_{2}(t)$ are non-exceptional, and
    $([\phi_{1}(t)],[\phi_{2}(t)])\in\mathrm{PNI}_{d};$
  • for $1\leq j\leq 2$, the morphism $\Psi\circ\phi_{j}:C\to\mathrm{MPoly}^{d}$ factors through $\Psi\circ\eta_{0}:\mathbb{A}^{d-1}\to\mathrm{MPoly}^{d}$ via some morphism $\psi_{j}:C\to\mathbb{A}^{d-1}$ over $\overline{\mathbb{Q}}$.
If $K$ is a number field such that $\phi_{1}$, $\phi_{2}$, and $C$ are defined over $K$, then one can verify that the base change of $C$, $\psi_{1}$, and $\psi_{2}$ in above reduction can also be defined over $K$. As all the properties we consider (multiplier spectrum, equivalence, exceptionalness etc.) do not change under conjugacy, we may replace $\phi_{j}$ by $\eta_{0}\circ\psi_{j}$ and assume that $\phi_{j}:C\to\mathrm{Poly}^{d}$ factors through $\eta_{0}:\mathbb{A}^{d-1}\to\mathrm{Poly}^{d}$ by $\psi_{j}$, for $1\leq j\leq 2$.
By Proposition 4.3, we obtain:
  • $m(\phi_{1})=m(\phi_{2})$ and $Z(\phi_{1})=Z(\phi_{2})$;
  • $Z^{\prime}(\phi_{1})=\#Z^{\prime}(\phi_{2})=+\infty$ and $\left(Z^{\prime}(\phi_{1})\Delta Z^{\prime}(\phi_{2})\right)<+\infty$.
For $1\leq j\leq 2$ and
$k\in A(\phi_{j})=\{0\leq l\leq d-2\colon c_{l}\circ\psi_{j}\text{ is not persistently }\phi_{j}\text{-preperiodic}\},$
set
$Z_{jk}:=\{t\in C(\mathbb{C})\colon c_{k}(\psi_{j}(t))\in\mathrm{Per}(\phi_{j}(t))\}=\{t\in C(\overline{\mathbb{Q}})\colon c_{k}(\psi_{j}(t))\in\mathrm{Per}(\phi_{j}(t))\},$
@@
where the last equality is because all the objects are defined over $\overline{\mathbb{Q}}$. By definition, $Z^{\prime}(\phi_{j})=\bigcup_{k\in A(\phi_{j})}Z_{jk}$ for $1\leq j\leq 2$. Since the intersection of $Z^{\prime}(\phi_{1})$ and $Z^{\prime}(\phi_{2})$ is infinite, we can choose $(k,l)\in A(\phi_{1})\times A(\phi_{2})$ such that $(Z_{1k}\cap Z_{2l})=+\infty$. After renumbering, assume $k=l=0$. Then for $1\leq j\leq 2$, the marked critical point 0 is not persistently $\phi_{j}$-preperiodic, hence $(\phi_{j},0)$ is an active dynamical pair. We have
$\{t\in C(\overline{\mathbb{Q}})\colon 0\in\mathrm{Per}(\phi_{1}(t))\cap\mathrm{Per}(\phi_{2}(t))\}=+\infty,$
hence the pairs $(\phi_{1},0)$ and $(\phi_{2},0)$ are entangled in the sense of Favre and Gauthier [FG22].
@@
Applying Theorem 4.1, there exist $N,M\in\mathbb{Z}_{>0}$ and families $R,\tau,\pi$ of polynomials of degree $\geq 1$ parametrized by $C$ such that
$\tau\circ\phi_{1}^{\circ N}=R\circ\tau\quad\text{and}\quad\pi\circ\phi_{2}^{\circ M}=R\circ\pi.$
Note that $d^{N}=\deg(\phi_{1}^{\circ N})=\deg(R)=\deg(\phi_{2}^{\circ M})=d^{M}$, so $N=M$.
@@
Fix $t\in C(\mathbb{C})$. Write $h_{1}=\phi_{1}(t)^{\circ N}$ and $h_{2}=\phi_{2}(t)^{\circ N}$. For simplicity, we abuse the notation and write $R,\tau,\pi$ for $R(t),\tau(t),\pi(t)$, respectively. Then
$\tau\circ h_{1}=R\circ\tau\quad\text{and}\quad\pi\circ h_{2}=R\circ\pi.$
Assume that $h_{1}$ and $h_{2}$ are not equivalent, i.e., the statement (1) of Theorem 1.10 does not hold for $N$ and the fixed $t\in C(\mathbb{C})$.
@@
Claim: There exist integers $k_{1},k_{2},l>0$, a non-constant polynomial $V(z)\in\mathbb{C}[z]\setminus\mathbb{C}$, and $\zeta\in\mathbb{C}$, with
$2\leq k:=\mathrm{lcm}(k_{1},k_{2})\leq d^{N},\quad\gcd(d,k)=1,\quad V(0)\neq 0,\quad\zeta^{k_{2}}=1,$
@@
where $k_{j}^{\prime}=k/k_{j}$ for $1\leq j\leq 2$.
We prove the Claim using results of Pakovich and elementary argument. Since $h_{1}\not\sim h_{2}$, we have $h_{j}\not\sim R$ for some $1\leq j\leq 2$. Without loss of generality, assume $h_{1}\not\sim R$. By Theorem 2.10, $R$ is a generalized Lattès map and there exist $\tau_{0},\tau_{1},\tilde{h}_{1}\in\mathbb{C}(z)\setminus\mathbb{C}$ satisfying:
  • $\tau=\tau_{0}\circ\tau_{1}$ and $k_{1}:=\deg(\tau_{0})\geq 2$;
@@
  • Both $R:\mathcal{O}_{2}^{\tau_{0}}\to\mathcal{O}_{2}^{\tau_{0}}$ and $\tilde{h}_{1}:\mathcal{O}_{1}^{\tau_{0}}\to\mathcal{O}_{1}^{\tau_{0}}$ are minimal holomorphic maps between orbifolds;